MINUTES FROM THE BOOSTER SUPERVISOR MEETING
Links: Minutes, PSB home page, PSB
PRESENT: M. Benedikt, A. Blas, M. Chanel, G. Cyvoct, G. Gräwer, C.
Supervisor Linac: R. Scrivens à G. Gräwer
Supervisor PSB: G. Cyvoct à E. Jensen
LINAC: (smooth running)
- Monday, 17. June: Tube in pre-amplifier of RF chain Tank 2 was replace, OK
- C. Hill took the opportunity of this intervention to have a closer look at
the source, which had too many breakdowns some time ago (not recently
though). He saw quite some dust, but did not have to intervene.
- Many reboots of several DSC's were necessary over the week (in particular
all RF DSC's which is annoying because their reboot stops the beam.)
- The coupled parameters which are used at injection (Injection
in LINC were (by mistake) made PPM. Corrected and OK now.
- On Friday, 14th June it was found on user TOF, that at ejection there was
still a high (> 2 kV) component of h=2, which in turn caused substantial
losses at PS injection. This has been corrected (the h=2 component of the RF
voltage is reduced just above the threshold of instability). These settings
were also copied to the parasitic TOF cycle, i.e. EASTA ring 2.
- Ejection septum BE.SMH: J.-M. Cravero had to adjust the active filter
again (the ACQ value was about 10 A too low).
- Vistar: the trouble with the Vistar is continuing. However, it seems
understood now that it is not a software bug, but rather a communication
problem (overloading the network). à suivre ...
- High intensity (in particular ISOGPS): Problem in ring 4 at or close to synchronization/ejection:
unstable longitudinally, which severely limits the intensity: To be looked
at! Today, 18. June, a similar behavior was also observed in ring 2!
- M. Chanel observed in this context that on ISOGPS, directly after
injection that there is a longitudinal oscillation at approx. 8 kHz
(4-pole?). In particular ring 1 is bad in this respect!
- This morning (18. June) M. George intervened on BT.BHZ10 in order to
equalize CCV and AQN. As a consequence we have to adjust all CCV's (PS
complained about bad trajectory). The values:
- towards PS: 356 A à
- towards ML: -375 A à
- Next BS meeting Tuesday 25-June-02 at 10.00 in the MCR Glassbox.
Minutes by Erk Jensen